OpenStudy (gabylovesyou):

on the graph of the equation 3x - 2y = -18, what is the value of the y-intercept?

OpenStudy (gabylovesyou):

@ganeshie8

OpenStudy (the_fizicx99):

9

OpenStudy (anonymous):

it is 6

OpenStudy (gabylovesyou):

how ?

OpenStudy (anonymous):

since move the x to the right divide by 3

hartnn (hartnn):

find it out : to get y-intercept, plug in x=0

OpenStudy (arabpride):

Begin by setting up the equation in slope-intercept form, aka y = mx + b to make your life a whole lot easier :3

hartnn (hartnn):

-2y =-18 getting y is just one step

OpenStudy (the_fizicx99):

Cute, just plot it on a coordinate grid..

OpenStudy (arabpride):

It would help if we gave her ONE way to solve it lel cx

OpenStudy (triciaal):

what is the value of the y-intercept? the y-intercept is the value of y when x = 0

OpenStudy (anonymous):

yes it would lol

OpenStudy (gabylovesyou):

ugh.. .-. ok so i plug in 0 ? 3*0 - 2y = -18

hartnn (hartnn):

yes, continue

hartnn (hartnn):

3*0 =0 divide throughout by -2

OpenStudy (gabylovesyou):

3 *0 - 2y = -18 0 - 2y = -18 correct ?

hartnn (hartnn):

yes

OpenStudy (triciaal):

yes

hartnn (hartnn):

y = -18/-2 = 18/2 = ..

OpenStudy (anonymous):

yupppp

OpenStudy (gabylovesyou):

so its 9

hartnn (hartnn):

correct :)

OpenStudy (gabylovesyou):

-18 / -2 = 9

OpenStudy (gabylovesyou):

Thank you.

hartnn (hartnn):

welcome ^_^